OpenStudy (anonymous):

Please help me solve Rob is 2 more than 3 times teds age. Four years ago rob was 5 times teds age. How old is rob?

OpenStudy (tkhunny):

Half the solution is in a good definition. Try this: R = Rob's Age Now. T = Ted's Age Now There you go. What's your plan for the other half the solution?

OpenStudy (anonymous):

Idk what to do of the second half

OpenStudy (anonymous):

r=2+3t

OpenStudy (anonymous):

Is that the first equation

OpenStudy (anonymous):

r-4=5t

OpenStudy (anonymous):

How do you solve these though

OpenStudy (tkhunny):

That's not good enough, natth88. You must learn to translate the words, given the clear definitions. "Rob is 2 more than 3 times teds age" R = 2 + 3*T Stare at that until you see it, then you do the next translation: "Four years ago rob was 5 times teds age. "